OpenStudy (karinamtz0213):

Please help!!Which of the following expressions are equivalent to (2+i)^2 where i is the imaginary number: I. (-2-i)^2 II. 3+4i III. 3 a. I only b. II only c. III only d. I and II only e. I and III only

OpenStudy (anonymous):

\[(2+i)^2=(2+i)(2+i)=(4-1)+(2+2)i=3+4i\]

OpenStudy (anonymous):

so II is good

OpenStudy (anonymous):

try and see if you get the same with \((-2-i)^2\)

OpenStudy (karinamtz0213):

okay thank you :)

OpenStudy (anonymous):

satellite can you help me http://openstudy.com/study#/updates/52044441e4b010d6e9b34027